OpenStudy (anonymous):

find an exact value. cos(-7pi/12)

OpenStudy (jdoe0001):

do you recall the sum trig identities? the ones you used in the previous one

OpenStudy (jdoe0001):

pictures isn't loading well anyhow \(\bf -\cfrac{7}{12}\implies \cfrac{\cancel{ 3 }}{\cancel{ 12 }}-\cfrac{\cancel{ 10 }}{\cancel{ 12 }}\implies \cfrac{1}{4}-\cfrac{5}{4}\qquad thus \\ \quad \\ -\cfrac{7\pi}{12}\implies\cfrac{\pi}{4}-\cfrac{5\pi}{4}\qquad thus \\ \quad \\ cos\left(-\frac{7\pi}{12}\right)\implies cos\left(\frac{\pi}{4}-\frac{5\pi}{4}\right)\)

OpenStudy (jdoe0001):

hmmm actaully lemme fix that

OpenStudy (jdoe0001):

\(\bf -\cfrac{7}{12}\implies \cfrac{\cancel{ 3 }}{\cancel{ 12 }}-\cfrac{\cancel{ 10 }}{\cancel{ 12 }}\implies \cfrac{1}{4}-\cfrac{5}{6}\qquad thus \\ \quad \\ -\cfrac{7\pi}{12}\implies\cfrac{\pi}{4}-\cfrac{5\pi}{6}\qquad thus \\ \quad \\ cos\left(-\frac{7\pi}{12}\right)\implies cos\left(\frac{\pi}{4}-\frac{5\pi}{6}\right)\) so use the sum trig identities, as shown in the previous one you did